can you please help me to answer thi question thank you
mang cesar is20 olderb than his son. in 5 years,he will be 3 times as old as his son.find their present age

Answer by dukesmart(20)

let x= present dad age; y=present son age
establis the equation as follows:
x-20=y .........(1)
x+5=3(y+5) .....(2)
take (1) subtract from (2) we have:
x+5-x+20 = 3y+15-y
2y= 25-15
2y = 10
y = 5
Present son age is 5 years old and Dad present age is x= x+20=5+20
x= 25 years old